OpenStudy (anonymous):

A school record in a certain race in 1960 was 3.8 minutes. In 1980 the school record in the same event was 3.65 minutes

OpenStudy (anonymous):

it would be b. from 1960 to 1980 the time decrease by 0.15 minutes, so slope would be −(0.15/20)=−0.0075 as b is the only equation with negative slope, it has to be the right one. Also you could just insert the numbers (x=year) if x=1960 then: −0.0075∗1960+18.5=−14.7+18.5=3.8 if x=1980 then −0.0075∗1980+18.5=−14.85+18.5=3.65